# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@ 3.1 Asymptotic Analysis
$2_{s}^{*}=\frac{2n}{n-2s},\quad\phi(x)=\frac{1-|x|^{2}}{2},\quad dV_{\mathbb{B}^{n}}=\phi(x)^{-n}\,dx\quad\text{on }\mathbb{B}^{n}.$
@@
$U(x):=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),$
@@
Let $0<s<\frac{n}{2}$. Then
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx=\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx+O(\varepsilon^{n})\quad{\rm as}\ \,\varepsilon\to 0,$
where the $O(\varepsilon^{n})$ constant depends only on $n,s$.
@@
$\int_{|y|\geq R}(1+|y|^{2})^{-n}\,dy\leq|\mathbb{S}^{n-1}|\int_{R}^{\infty}r^{n-1}r^{-2n}\,dr=\frac{|\mathbb{S}^{n-1}|}{n}\,R^{-n}.$
@@
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{p}\,dx=\int_{\mathbb{R}^{n}}|U_{\varepsilon}|^{p}\,dx+O(\varepsilon^{n})=\int_{\mathbb{R}^{n}}|U|^{p}\,dx+O(\varepsilon^{n}).$
@@
Let $0<s<\frac{n}{2}$. Then as $\varepsilon\to 0^{+}$
@@
$\int_{B_{2\delta}\setminus B_{\delta}}|\eta U_{\varepsilon}|^{2}\,dx=O(\varepsilon^{\,n-2s}).$ (3.4)
@@
$\int_{B_{2\delta/\varepsilon}}|y|^{2}(1+|y|^{2})^{-(n-2s)}\,dy=\begin{cases}O(1),&n>4s+2,\\[2.84526pt] O\bigl(|\log\varepsilon|\bigr),&n=4s+2,\\[2.84526pt] O\bigl(\varepsilon^{-(4s-n+2)}\bigr),&n<4s+2,\end{cases}\qquad\varepsilon\to 0.$ (3.7)
@@
$\int_{\mathbb{R}^{n}}|x|^{2}|w_{\varepsilon}(x)|^{2}\,dx=\begin{cases}O(\varepsilon^{2s+2}),&n>4s+2\\[2.84526pt] O\bigl(\varepsilon^{2s+2}|\log\varepsilon|\bigr),&n=4s+2\\[2.84526pt] O(\varepsilon^{n-2s}),&n<4s+2\end{cases}\qquad{\rm as}\ \,\varepsilon\to 0.$
@@
Let $0<s<\frac{n}{2}$ and set $a=\frac{n-2s}{2}$. Define
$U(x)=(1+|x|^{2})^{-a},\qquad U_{\varepsilon}(x)=\varepsilon^{-a}\,U\!\left(\frac{x}{\varepsilon}\right).$
Then for every multi-index $\alpha\in\mathbb{N}_{0}^{n}$ and every $\delta>0$ there exists $C_{\alpha,\delta}>0$ such that for all $\varepsilon\in(0,1)$ and all $x\in\mathbb{R}^{n}$ with $|x|\geq\delta$,
@@
Step 1: Scaling invariance of the main term. Recall
$U_{\varepsilon}(x)=\varepsilon^{-\frac{n-2s}{2}}\,U\!\left(\frac{x}{\varepsilon}\right).$
We claim that
@@
Proof of Theorem 1.3: Let
$U(x):=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),$
@@
$B_{\varepsilon}:=\Bigl(\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{{\mathbb{B}}^{n}}\Bigr)^{\!\frac{2}{2_{s}^{*}}}=\Bigl(\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}},$
by Proposition 3.2, we have $B_{\varepsilon}=B_{0}+O(\varepsilon^{n})$ with
$B_{0}:=\Bigl(\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}>0,$
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}\;=\;S_{n,s}\;+\;O(\varepsilon^{\,n-2s}).$ (3.23)
@@
$S_{n,s}\;\leq\;\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}\;+\;\lambda\,\frac{\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}.$ (3.24)
@@
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}+O(\varepsilon^{n-2s})\qquad\text{with }\widetilde{c}_{n,s}>0,$
@@
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}|\log\varepsilon|+O(\varepsilon^{2s}),\qquad\widetilde{c}_{n,s}>0,$
@@
Combining this estimate with [24, Theorem 1.8], we can only conclude that
@@
(i) Applying [24, Lemma 5.2] with $\gamma=s$ and $\lambda=0$, we obtain
@@
$M_{s}(\beta):=\Bigl|\Gamma\Bigl(s+\tfrac{1}{2}+i\beta\Bigr)\Bigr|^{2},\qquad\beta\geq 0.$
@@
$\lim_{|\lambda|\to\infty}\,\bigl|\Gamma(a+i\lambda)\bigr|\,e^{\frac{\pi}{2}|\lambda|}\,|\lambda|^{\frac{1}{2}-a}=\sqrt{2\pi},$
@@
$\mathcal{P}_{s}=\widetilde{\mathcal{P}}_{s}+\frac{\sin(s\pi)}{\pi}\,M_{s}(A),\ s\in\left(0,\frac{n}{2}\right)\setminus\mathbb{N}$ (3.28)
@@ 4.1 Energy Spaces
and for exponents
$1<p\leq 2_{s}^{*}-1,\qquad 2_{s}^{*}:=\frac{2n}{n-2s}.$
Since the bottom spectrum is
@@
Consequently, by [19, Theorem E.8], we obtain the sharp fractional Poincaré inequality
@@
$\langle u,v\rangle_{\lambda,s,\sim}:=\langle(\widetilde{\mathcal{P}}_{s}-\lambda)u,v\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}v\ (\widetilde{\mathcal{P}}_{s}-\lambda)u,dV_{\mathbb{H}^{n}},\qquad\|u\|_{\lambda,s,\sim}^{2}:=\langle u,u\rangle_{\lambda,s,\sim}={\widetilde{\mathcal{E}}}_{\lambda,s}(u).$
By [24, Theorem 1.3 and Theorem 1.8], for every $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s}$ defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$, and for every $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s,\sim}$ also defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$
@@
Let $s\in(0,\frac{n}{2})$. $\widetilde{m}_{s}(\beta)$ is defined in (2.8). Then there exist constants $c_{1},c_{2}>0$ (depending only on $s$) such that for all $\beta\in\mathbb{R}$,
@@
Proof: First note that $\widetilde{m}_{s}$ is even and strictly positive on $\mathbb{R}$, since $\Gamma$ has no zeros and $\Gamma(\tfrac{1}{2}+i\beta)\neq 0$ for all $\beta\in\mathbb{R}$. We use the uniform Stirling estimate in vertical strips: for each fixed $\sigma>0$, there exist $T\geq 1$ and constants $C_{\pm}=C_{\pm}(\sigma)>0$ such that for all $\sigma^{\prime}\in[\sigma,\sigma+1]$ and all $|t|\geq T$,
$C_{-}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}\ \leq\ |\Gamma(\sigma^{\prime}+it)|\ \leq\ C_{+}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}.$ (4.2)
@@
$\frac{\bigl|\Gamma(s+\frac{1}{2}+i\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+i\beta)\bigr|}\sim|\beta|^{s},$
@@
$\widetilde{m}_{s}(\beta)=\Bigl(\frac{\bigl|\Gamma(s+\frac{1}{2}+i\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+i\beta)\bigr|}\Bigr)^{\!2}\sim|\beta|^{2s},\qquad|\beta|\geq T.$
@@
$1+m_{s}(\beta)^{2}\ \sim\ 1+(\beta^{2}+1)^{2s},\qquad 1+\beta\in\mathbb{R},$ (4.4)
@@
Let $s\in(0,\frac{n}{2})\setminus\mathbb{N},$ for every $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, the quadratic form
$\|u\|_{\lambda,s}^{2}:=\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})},\quad u\in C_{c}^{\infty}(\mathbb{H}^{n})$
extends to a norm on $H^{s}(\mathbb{H}^{n})$, which is equivalent to the standard $H^{s}$-norm.
Proof: By Plancherel formula,
@@
To bound $1+m_{s}(\beta)$ from above by $m_{s}(\beta)-\lambda$, set
@@
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\ \|\cdot\|_{\lambda,s}},\quad\lambda<\lambda_{0,s}^{\mathrm{conf}}.$
@@
Throughout, we write $\|u\|_{p+1}:=\|u\|_{L^{p+1}(\mathbb{H}^{n})}$ for brevity. It is standard to work on the Nehari manifold
$\mathcal{N}_{\lambda,s}:=\Bigl\{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}:\mathcal{E}_{\lambda,s}(u)=\|u\|_{p+1}^{p+1}\Bigr\},$
@@
Let $\mathcal{T}$ is an isometry on $\mathbb{H}^{n}$ and define
@@
$\mathcal{P}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\mathcal{P}_{s}u),\quad\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\widetilde{\mathcal{P}}_{s}u)$
@@
$\langle U_{\mathcal{T}}u,\mathcal{P}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\mathcal{P}_{s}u\rangle_{2},\quad\langle U_{\mathcal{T}}u,\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\widetilde{\mathcal{P}}_{s}u\rangle_{2}$
Proof: Since $\mathcal{T}$ is an isometry, and $dV_{\mathbb{H}^{n}}$ is invariant under $\mathcal{T}$, hence $U_{\mathcal{T}}$ is unitary on $L^{2}(\mathbb{H}^{n})$ and preserves $L^{p}$ norms by change of variables. Moreover, the Laplace–Beltrami operator is invariant under isometries, i.e. $\Delta_{\mathbb{H}^{n}}(u\circ\mathcal{T})=(\Delta_{\mathbb{H}^{n}}u)\circ\mathcal{T}$. Therefore $U_{\mathcal{T}}^{-1}(-\Delta_{\mathbb{H}^{n}})U_{\mathcal{T}}=-\Delta_{\mathbb{H}^{n}}$. Since on $\mathbb{H}^{n}$, $P_{s}$ and $\widetilde{P}_{s}$ are given by spectral calculus as a Borel function of
$A:=-\Delta_{\mathbb{H}^{n}}-\rho^{2},$
@@
$\mathcal{P}_{s}(u\circ\mathcal{T})=(\mathcal{P}_{s}u)\circ\mathcal{T},\quad\widetilde{\mathcal{P}}_{s}(u\circ\mathcal{T})=(\widetilde{\mathcal{P}}_{s}u)\circ\mathcal{T}$

### v2

@@ 3.1 Asymptotic Analysis
$\phi(x)=\frac{1-|x|^{2}}{2},\quad dV_{\mathbb{B}^{n}}=\phi(x)^{-n}\,dx\quad\text{on }\mathbb{B}^{n}.$
@@
$U(x)=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x)=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),\quad\forall x\in\mathbb{R}^{n}$
@@
Let $s\in(0,\frac{n}{2})$, then
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx=\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx+O(\varepsilon^{n})\quad{\rm as}\ \,\varepsilon\to 0^{+},$
where $O(\varepsilon^{n})=\varepsilon^{n}O(1)$ constant depends only on $n,s$.
@@
$\int_{|y|\geq R}(1+|y|^{2})^{-n}\,dy\leq|\mathbb{S}^{n-1}|\int_{R}^{\infty}r^{n-1}r^{-2n}\,dr=\frac{|\mathbb{S}^{n-1}|}{n}\,R^{-n},$
@@
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{p}\,dx=\int_{\mathbb{R}^{n}}|U_{\varepsilon}|^{p}\,dx+O(\varepsilon^{n})=\int_{\mathbb{R}^{n}}|U|^{p}\,dx+O(\varepsilon^{n})\quad{\rm as}\ \,\varepsilon\rightarrow 0^{+}.$
We next derive the $L^{2}$-mass estimate in the hyperbolic space. In contrast to the Euclidean case, the asymptotic behavior is influenced by the conformal weight and must be analyzed with respect to the hyperbolic volume element.
@@
Let $0<s<\frac{n}{2}$, then as $\varepsilon\to 0^{+}$
@@
$\int_{B_{2\delta}\setminus B_{\delta}}|\eta U_{\varepsilon}|^{2}\,dx=O(\varepsilon^{\,n-2s})\quad{\rm as}\ \,\varepsilon\rightarrow 0^{+}.$ (3.4)
@@
$\int_{B_{2\delta/\varepsilon}}|y|^{2}(1+|y|^{2})^{-(n-2s)}\,dy=\begin{cases}O(1),&n>4s+2,\\[2.84526pt] O\bigl(|\log\varepsilon|\bigr),&n=4s+2,\\[2.84526pt] O\bigl(\varepsilon^{-(4s-n+2)}\bigr),&n<4s+2,\end{cases}\qquad{\rm as}\ \,\varepsilon\to 0^{+}.$ (3.7)
@@
$\int_{\mathbb{R}^{n}}|x|^{2}|w_{\varepsilon}(x)|^{2}\,dx=\begin{cases}O(\varepsilon^{2s+2}),&n>4s+2\\[2.84526pt] O\bigl(\varepsilon^{2s+2}|\log\varepsilon|\bigr),&n=4s+2\\[2.84526pt] O(\varepsilon^{n-2s}),&n<4s+2\end{cases}\qquad{\rm as}\ \,\varepsilon\to 0^{+}.$
@@
The next derivative estimate will be used to control the error terms coming from the cut-off region away from the concentration point, which is vital in proof of Proposition 3.5.
@@
Let $0<s<\frac{n}{2}$ and set $a=\frac{n-2s}{2}$, then for every multi-index $\alpha\in\mathbb{N}_{0}^{n}$ and every $\delta>0$ there exists $C_{\alpha,\delta}>0$ such that for all $\varepsilon\in(0,1)$ and all $x\in\mathbb{R}^{n}$ with $|x|\geq\delta$,
@@
Step 1: Scaling invariance of the main term. We claim that
@@
With the asymptotic estimates established above for the cut-off bubble, we are now in a position to prove Theorem 1.3 and Theorem 1.4.
Proof of Theorem 1.3: Recall
$U(x)=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x)=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),$
@@
$L_{\varepsilon}:=\Bigl(\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{{\mathbb{B}}^{n}}\Bigr)^{\!\frac{2}{2_{s}^{*}}}=\Bigl(\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}},$
by Proposition 3.2, we have $L_{\varepsilon}=L_{0}+O(\varepsilon^{n})$ with
$L_{0}:=\Bigl(\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}>0,$
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{L_{\varepsilon}}\;=\;S_{n,s}\;+\;O(\varepsilon^{\,n-2s}).$ (3.23)
@@
$S_{n,s}\;\leq\;\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{L_{\varepsilon}}\;+\;\lambda\,\frac{\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{L_{\varepsilon}}.$ (3.24)
@@
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{L_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}+O(\varepsilon^{n-2s})\qquad\text{with }\widetilde{c}_{n,s}>0,$
@@
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{L_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}|\log\varepsilon|+O(\varepsilon^{2s}),\qquad\widetilde{c}_{n,s}>0,$
@@
Combining this estimate with [29, Theorem 1.8], we can only conclude that
@@
(i) Applying [29, Lemma 5.2] with $\gamma=s$ and $\lambda=0$, we obtain
@@
$M_{s}(\beta):=\Bigl|\Gamma\Bigl(s+\tfrac{1}{2}+{\bf i}\beta\Bigr)\Bigr|^{2},\quad\beta\geq 0.$
@@
$\lim_{|\lambda|\to\infty}\,\bigl|\Gamma(a+{\bf i}\lambda)\bigr|\,e^{\frac{\pi}{2}|\lambda|}\,|\lambda|^{\frac{1}{2}-a}=\sqrt{2\pi},$
@@
$\mathcal{P}_{s}=\widetilde{\mathcal{P}}_{s}+\frac{\sin(s\pi)}{\pi}\,M_{s}(A),\ s\in\left(0,\frac{n}{2}\right)\setminus\mathbb{N},$ (3.28)
@@ 4  Fractional Brezis–Nirenberg Problems
In this section, we study the attainability of the fractional Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, and, as an application, we establish existence results for the corresponding fractional Brezis–Nirenberg problems.
We first introduce the natural functional settings adapted to these operators. To handle the nonlocal difficulties—in particular, the lack of direct integration-by-parts identities and pointwise formulas, we develop suitable fractional energy estimates via the pseudodifferential operator framework on manifolds. These estimates play a central role in the attainability analysis.
We then show that once $\lambda$ exceeds the spectral bottom of the corresponding operator, the associated Poincaré–Sobolev level drops to -$\infty$. A key ingredient is the off-diagonal exponential decay estimate (see Proposition 4.10). To prove this proposition, we exploit the fact that for fractional operators the most accessible information is encoded on the Fourier side through explicit multipliers, whereas direct control in physical space is less immediate. We therefore combine the Schwartz kernel theorem, which provides a distributional kernel representation, with the Harish–Chandra asymptotic expansion to derive the required off-diagonal decay.
@@ 4.1 Energy Spaces
for $1<p\leq 2_{s}^{*}-1.$ From the bottom spectrum
@@
and by [23, Theorem E.8], we obtain the sharp fractional Poincaré inequality
@@
$\langle u,v\rangle_{\lambda,s,\sim}:=\langle(\widetilde{\mathcal{P}}_{s}-\lambda)u,v\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}v\ (\widetilde{\mathcal{P}}_{s}-\lambda)u\,dV_{\mathbb{H}^{n}},\qquad\|u\|_{\lambda,s,\sim}^{2}:=\langle u,u\rangle_{\lambda,s,\sim}={\widetilde{\mathcal{E}}}_{\lambda,s}(u).$
By [29, Theorem 1.3 and Theorem 1.8], for every $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s}$ defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$, and for every $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s,\sim}$ also defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$
The next two lemmas show that these norms are equivalent to the standard spectral fractional Sobolev norm.
@@
Let $s\in(0,\frac{n}{2})$ and $\widetilde{m}_{s}(\beta)$ be defined in (2.9), then there exist constants $c_{1},c_{2}>0$ (depending only on $s$) such that for all $\beta\in\mathbb{R}$,
@@
Proof: First note that $\widetilde{m}_{s}$ is even and strictly positive on $\mathbb{R}$, since $\Gamma$ has no zeros and $\Gamma(\tfrac{1}{2}+{\bf i}\beta)\neq 0$ for all $\beta\in\mathbb{R}$. We use the uniform Stirling estimate in vertical strips: for each fixed $\sigma>0$, there exist $T\geq 1$ and constants $C_{\pm}=C_{\pm}(\sigma)>0$ such that for all $\sigma^{\prime}\in[\sigma,\sigma+1]$ and all $|t|\geq T$,
$C_{-}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}\ \leq\ |\Gamma(\sigma^{\prime}+{\bf i}t)|\ \leq\ C_{+}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}.$ (4.2)
@@
$\frac{\bigl|\Gamma(s+\frac{1}{2}+{\bf i}\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+{\bf i}\beta)\bigr|}\sim|\beta|^{s},$
@@
$\widetilde{m}_{s}(\beta)=\Bigl(\frac{\bigl|\Gamma(s+\frac{1}{2}+{\bf i}\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+{\bf i}\beta)\bigr|}\Bigr)^{\!2}\sim|\beta|^{2s},\qquad|\beta|\geq T.$
@@
$1+m_{s}(\beta)^{2}\ \sim\ 1+(\beta^{2}+1)^{2s}\quad{\rm for}\ \,\beta\in\mathbb{R},$ (4.4)
@@
Let $s\in(0,\frac{n}{2})\setminus\mathbb{N},$ then for every $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, the quadratic form
$\|u\|_{\lambda,s}:=\sqrt{\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})}},\quad\forall u\in C_{c}^{\infty}(\mathbb{H}^{n})$
can extend to a norm on $H^{s}(\mathbb{H}^{n})$ and it is equivalent to the standard $H^{s}$-norm.
Proof: By Plancherel formula, one has
@@
To get the bound $1+m_{s}(\beta)$ from above, set
@@
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\ \|\cdot\|_{\lambda,s}}\quad{\rm for}\ \,\lambda<\lambda_{0,s}^{\mathrm{conf}}.$
@@
Throughout, we write $\|u\|_{q}:=\|u\|_{L^{q}(\mathbb{H}^{n})}$ for brevity. It is standard to work on the Nehari manifold
$\mathcal{N}_{\lambda,s}:=\Bigl\{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}:\mathcal{E}_{\lambda,s}(u)=\|u\|_{p+1}^{p+1}\Bigr\}$
@@
The next lemma records the isometry invariance of $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, and in particular the invariance of their associated quadratic forms.
@@
Let $\mathcal{T}$ be an isometry on $\mathbb{H}^{n}$ and define
@@
$\mathcal{P}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\mathcal{P}_{s}u),\quad\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\widetilde{\mathcal{P}}_{s}u).$
@@
$\langle U_{\mathcal{T}}u,\mathcal{P}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\mathcal{P}_{s}u\rangle_{2},\quad\langle U_{\mathcal{T}}u,\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\widetilde{\mathcal{P}}_{s}u\rangle_{2}.$
Proof: Since $\mathcal{T}$ is an isometry, and $dV_{\mathbb{H}^{n}}$ is invariant under $\mathcal{T}$, hence $U_{\mathcal{T}}$ is unitary on $L^{2}(\mathbb{H}^{n})$ and preserves $L^{p}$ norms by change of variables. Moreover, the Laplace–Beltrami operator is invariant under isometries, i.e. $\Delta_{\mathbb{H}^{n}}(u\circ\mathcal{T})=(\Delta_{\mathbb{H}^{n}}u)\circ\mathcal{T}$. Therefore $U_{\mathcal{T}}^{-1}(-\Delta_{\mathbb{H}^{n}})U_{\mathcal{T}}=-\Delta_{\mathbb{H}^{n}}$. Since on $\mathbb{H}^{n}$, $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ are given by spectral calculus as a Borel function of
$\mathcal{A}^{2}:=-\Delta_{\mathbb{H}^{n}}-\rho^{2},$
@@
$\mathcal{P}_{s}(u\circ\mathcal{T})=(\mathcal{P}_{s}u)\circ\mathcal{T},\quad\widetilde{\mathcal{P}}_{s}(u\circ\mathcal{T})=(\widetilde{\mathcal{P}}_{s}u)\circ\mathcal{T}.$